lolokrat:

tutor told me #7-#13 were incorrect. can someone walk me through the steps of how to fix them?

dude:

#7 <4 and <5 are alternate interior, they are the same measure

dude:

#8,9 will be resolved once 7 is fixed Just make sure to re-do them

lolokrat:

for #7 arent i trying to find x?? how do i do that?

dude:

You set them equal to 180 Its supposed to be \(4x+15=6x-25\)

lolokrat:

wait do i do 4x+15=6x-25 ?

dude:

Right

lolokrat:

yea

lolokrat:

4x+15=6x-25 2x+15=-25 2x=-40 x=-20 is this right?

dude:

Not quite \(4x+15=6x-25\\ \color{red}{2x}+15=-25\\ 2x=-40\\ x=-20\) You got the wrong sign here 4-6 is -2 not 2

lolokrat:

oh 4x+15=6x−25 -2x+15=−25 -2x=−40 x=20

dude:

Yes

dude:

That is your answer to #7

lolokrat:

sorry, it takes me a little while to solve

dude:

For #8 Substitute 20 in for x \(4x+15\) Its okay, take your time

lolokrat:

4x+15 4(20)+15 80+15 95 <4=95º

dude:

Yeah, assuming that the measures are right You do have a point in saying that <1 is congruent to <4, but the teacher said it was wrong no?

lolokrat:

he just told me that #7-13 were incorrect

dude:

Okay then it was wrong, 95º is your answer

dude:

For #9 Substitute 20 in for x \(6x-25\)

lolokrat:

ok. 6x-25 6(20)-25 120-25 95 <5=95º

dude:

Yeh

lolokrat:

this may be a dumb question, but why are we substituting x with 20?

lolokrat:

oh from #7

dude:

Yes

lolokrat:

how would i do #10?

dude:

It looks right, I am not sure what went wrong hmm They are same side exterior which are supplementary

lolokrat:

yeah i had a friend tell me too that they thought that i was right. im not sure.

dude:

Maybe ask your tutor to re-check?

lolokrat:

ok. should we move onto #11, 12, and 13 with <7 being 50?

dude:

For #11 They also look right They are supplementary.. hmm #12 and 13 are also right because of this lolwut

lowkey:

What a good tutor, hunting out your mistakes rather than your successes.

dude:

Oh 13 re-check math

lolokrat:

@dude sorry, i fell asleep on you there. would <3=94.5?

lolokrat:

oh wait sorry i did 12

dude:

Uh \(3(21.8)+20\)

dude:

Yeah

lolokrat:

<3=85.4

dude:

Yeah